OpenStudy (anonymous):

Given theta=π/6, which of the following shows tan theta cot theta=1?

OpenStudy (campbell_st):

cot = 1/tan so the problem is \[\tan(\theta)\times \frac{1}{\tan(\theta)} = 1\] so the angle measure doesn't matter

OpenStudy (anonymous):